Sunland Corporation had a net income for the current year ending December 31, 2021 of $1,235,500. Throughout 2021 the following items were outstanding:

386,500 common shares
19,500 Class A $3 cumulative preferred shares that were convertible to common shares at a rate of 1:1
52,000 Class B $4 non-cumulative preferred shares that were convertible at a rate of one common share for every two preferred shares.
$504,000, 8% bonds that were convertible to 14,000 common shares
$297,000, 10% bonds convertible to 12,000 common shares


No dividends were declared or paid in 2021. Sunland’s tax rate is 25%.

1.Calculate the income effect of the dividends on Class A preferred shares.

2.Calculate the income effect of the dividends on Class B preferred shares.

3.Calculate basic earnings per share. (For simplicity, ignore the requirement to record the debt and equity portions of the convertible bond separately). (Round answer to 2 decimal places, e.g. 15.25.)

4.Calculate the after-tax interest paid on the 8% bonds.

5.Determine an incremental per share effect for 8% bonds. (Round earnings per share to 2 decimal places, e.g. 15.25.)

Potentially dilutive security Incremental
Numerator Effect
Incremental
Denominator Effect
EPS
8% Bonds $ $

6.Calculate the after-tax interest paid on the 10% bonds.

7.

Determine an incremental per share effect for 10% bonds. (Round earnings per share to 2 decimal places, e.g. 15.25.)

Potentially dilutive security Incremental
Numerator Effect
Incremental
Denominator Effect
EPS
10% Bonds $ $

8.Determine an incremental per share effect for Class A preferred shares. (Round earnings per share to 2 decimal places, e.g. 15.25.)

Potentially dilutive security Incremental
Numerator Effect
Incremental
Denominator Effect
EPS
Class A preferred shares $ $

9.Determine an incremental per share effect for Class B preferred shares. (Round earnings per share to 2 decimal places, e.g. 15.25.)

Potentially dilutive security Incremental
Numerator Effect
Incremental
Denominator Effect
EPS
Class B Preferred shares $ $

10.Rank the potentially dilutive securities from most dilutive to least dilutive.

8% bonds                                                                       Rank 1Rank 2Rank 3Rank 4Anti-dilutive
10% bonds                                                                       Rank 1Rank 2Rank 3Rank 4Anti-dilutive
Class A preferred shares                                                                       Rank 1Rank 2Rank 3Rank 4Anti-dilutive
Class B preferred shares                                                                       Rank 1Rank 2Rank 3Rank 4Anti-dilutive

11.Calculate diluted earnings per share. (Round calculations of EPS to 3 decimal places, e.g. $3.545 and provide final answer to 2 decimal places, e.g. 15.25.)

Numerator Denominator EPS
Basic $ $
                                                                      Class A preferred sharesClass B preferred shares8% bonds10% bonds
Sub Total
                                                                      Class A preferred sharesClass B preferred shares8% bonds10% bonds
Sub Total
                                                                      Class A preferred sharesClass B preferred shares8% bonds10% bonds
$ $
Diluted EPS

: P : qiii ? sm : i ?* ܙܬܙܬܐ 3 Que un 1 {Thank van ܗ ܀ ܕ: ܕ. ܗܬܝ܀. ܬܗܝ. ܀ ܐ 1 Trans % A4va ܪ ܙܕܕ T : : : : r - w ܙܓ݁ܙ. ܗܢ
Answer #1

Solution 1 : Dividend on Class A preferred shares = $19500*3 = $58500

Solution 2 : Dividend on Class B preferred Shares = $52000*4 = $208000

Solution 3 Calculation of Basic Earnings per share

                                $1,235,500 - $58500-$208000/386500 shares = $2.51 per share

Solution 4 : After tax interest on 8% bonds

                                = $504000*8% * (1-.25) = $30240